OpenStudy (anonymous):

How can we get dy/dx of sqrt(x+y)=1+x^2y^2 by using implicit differentiation?

OpenStudy (amistre64):

by using the rules of derivatives that were taught during explicits

OpenStudy (amistre64):

what would you say is the derivative of sqrt(u) ?

OpenStudy (wikiemol):

All you have to do is apply the chain rule. you know the derivative of y = dy/dx so anytime you are dealing with a derivative of y you multiply the derivative of the y related term by dy/dx. so for instance the derivative of xy^2 = y^2 + 2xy(dy/dx). Then, all you have to do is solve for dy/dx in terms of x and y.

OpenStudy (amistre64):

nice and succinct

OpenStudy (anonymous):

I see, thanks. After reading the textbook, I am still confused about some concept. For example, d/dx (7x+y)=7+dy/dx ? why should not equal 7+y*dy/dx?

OpenStudy (wikiemol):

well the derivative of a term such as a*u where a is a constant number is just a. so in this case we have the derivative of 1*u = 1 but we have to multiply it by the derivative of the inside which is dy/dx so we end up with d/dx (1*y) = 1*dy/dx or just dy/dx

OpenStudy (amistre64):

my method is to ignore that this is spose to be wrtx and just take derivative like normal D[7x+y] D[x] + D[7y] x' + 7 D[y] x' + 7y' now since x' wrtx = dx/dx = 1 ... 1 + 7y'

OpenStudy (amistre64):

\[\sqrt{x+y}=1+x^2y^2\] \[D[\sqrt{x+y}]=D[1+x^2y^2]\] \[\frac{D[x+y]}{2\sqrt{x+y}}=D[1]+D[x^2y^2]\] \[\frac{D[x]+D[y]}{2\sqrt{x+y}}=0+D[x^2]y^2+x^2D[y^2]\] \[\frac{x'+y'}{2\sqrt{x+y}}=2x~x'~y^2+x^2~2y~y'\]\[x'=1\] \[\frac{1+y'}{2\sqrt{x+y}}=2xy^2+2x^2y~y'\] and algebra it into y' = .....

OpenStudy (anonymous):

I see, thank you so much.